Calculate the relative molecular mass of the following gases if:
i. 0.4g of oxygen at stp of 280cm
1
Expert's answer
2021-07-08T04:54:39-0400

Solution:

Convert cm3 to L:

(280 cm3) × (0.001 L / 1 cm3) = 0.28 L


At STP, one mole of any gas occupies a volume of 22.4 L.

Thus 0.28 L of gas correspond to:

(0.28 L) × (1 mol / 22.4 L) = 0.0125 mol


Moles = Mass / Molecular mass

Thus,

Molecular mass of gas = Mass of gas / Moles of gas

Molecular mass of gas = (0.4 g) / (0.0125 mol) = 32 g/mol

Molecular mass of gas = 32 g/mol


Answer: The relative molecular mass of the gas is 32 g/mol